The reaction that takes place in the reusable solid rocket booster for the space shuttle is shown by the following balanced equation: 3 Al (s) + 3 NH4ClO4 (s) 6 Al2O3 (s) + AlCl3 (s) + 3 NO (g) + 6 H2O (g)

What volume of NO gas is formed from the reaction of 100 kg of Al (s) if the reaction is run at 1.03 atm and 300EC? What is the mass, in kg, of water (MM = 18.015 g/mol) produced?
Answer #1

Balanced equation:
3 Al(s) + 3 NH4ClO4(s) ====> Al2O3(s) + AlCl3(s) + 3 NO(g) + 6 H2O(g)

Mass of Al = 100 Kg = 100 x 10^3 g  

Molar mass of Al = 26.98 g/mol

Moles of Al = 100 x 10^3 g / 82.079 g/mol =  3706.238 Moles

Moles of NO produced   = 3706.238 Moles

According to the balanced equation one mole of Al will produce one mole of NO.

Moles of H2O produced   =7412.47 Moles

According to the balanced equation one mole of Al will produce two mole of H2O.

Mass of H2O produced = 133537.8 g = 133.537 Kg   

using ideal gas equation let us calcualte the volume of NO

PV= nRT

P = Pressure in atm                       V= Volume in Liter

n = no of moles               R = 0.08206 L atm K-1 Mol-1

T = Temperature in Kelvin = 300 Deg Cel = 573 K

Actually given in 300EC. So I took 300 Deg cel.

V = nRT / P  

V = 3706.238 Mole x 0.08206 L atm K-1 Mol-1 x 573 K / 1.03 atm  

V = 1,69,192.93 Liter

Hence the volume of NO =   1,69,192.93 Liter

Balanced equation: 3 Al(s) + 3 NH4C104($) = Al2O3(s) + AIC13(s) + 3 NO(g) + 6H2O(g) Reaction stoichiometry Limiting reagent C